Given that ellipse has an equation of 9x²+16y²=144.

Find(I)the x and y intercept of the graph of the equation

(II)the coordinates of the foci

(III)the length of the major and minor axis


1
Expert's answer
2021-02-01T15:04:32-0500

i) X intercept of the graph of the equation

"9x^2+16(0)^2=144\\x=\pm4"

Y intercept of the graph of the equation


"9(0)^2+16y^2=144\\y=\pm3"

ii) The coordinates of the foci


"(\pm\sqrt{4^2-3^2},0)=(\pm\sqrt{7},0)"

iii)


"L=2a=2(4)=8"

"l=2b=2(3)=6"